# Semiclassical Gravity Beyond General Relativity: Insights from Torsion

### v1

@@ 1 Introduction
Importantly, the empirical success of GR does not rule out the presence of torsion. In fact, Einstein–Cartan theory reproduces all experimentally tested predictions of GR [12]. Whas is more, within the framework of the Standard Model, possible signatures of torsion can be sought as new interactions involving polarized Dirac spinors [13].
@@
We organize the paper as follows. In Sec. 2, we provide a brief overview of semiclassical gravity and introduce some technical aspects. Section 3 presents the Einstein–Cartan theory, and in Sec. 4, we derive the corresponding Hadamard bi-parametrix. In Sec. 5, which constitutes the core of the paper, we introduce the axiomatic framework employed in the renormalization scheme, identify the resulting ambiguities, and review the conformal anomaly. Finally, Sec. 6 contains our conclusions. Some useful expressions are collected in the Appendix.
@@ 2 Preliminaries
We begin by discussing the left-hand side of Eq. (1). The metric curvature tensor, ${\mathring{R}_{abc}}^{\ \ \ d}$, is constructed using the torsion-free and metric-compatible derivative operator $\mathring{\nabla}_{a}$. These two properties mean, respectively, that
@@
The only nontrivial trace of ${\mathring{R}_{abc}}^{\ \ \ d}$ is $\mathring{R}_{ab}={\mathring{R}_{acb}}^{\ \ \ c}$, which defines the Ricci tensor and satisfies $\mathring{R}_{ab}=\mathring{R}_{ba}$. Moreover, the trace of the Ricci tensor is the Ricci scalar, $\mathring{R}=g^{ab}\mathring{R}_{ab}$. Finally, the Einstein tensor is given by
@@
As we shall see, this issue is resolved by acting with the energy–momentum operator on the Green function, $G(x,x^{\prime})$, and then taking the coincidence limit $x^{\prime}\rightarrow x$. This evaluation is divergent, and a renormalization procedure is required. To this end, Hadamard states are employed; these are states whose singular structure mirrors that of quantum fields in flat spacetime [30, 31, 32]. This allows the subtraction of the singular part (an alternative approach is discussed in Ref. [33]). Importantly, in semiclassical GR, an axiomatic framework exists [22] that leads to a well-defined expectation value for the energy–momentum tensor.
Our goal is to study semiclassical gravity in the Einstein–Cartan theory, where spacetime torsion is nontrivial. The procedure we employ closely follows that of Décanini and Folacci [32]. The most notable difference compared with GR is that the energy–momentum tensor is no longer divergence-free, a property that is repeatedly used in Ref. [32]. Instead, in Einstein–Cartan theory, its divergence satisfies a specific identity that also involves torsion and the spin–density tensor. We now turn to study some classical aspects of this theory.
@@ 3.1 Kinematics
We consider a nontrivial torsion tensor $T^{c}_{\ ab}$, defined by
$(\nabla_{a}\nabla_{b}-\nabla_{b}\nabla_{a})=-T^{c}_{\ ab}\nabla_{c}f,$ (6)
for any function $f$. By definition, $T^{c}_{\ ab}=T^{c}_{\ [ab]}$. Here, $\nabla_{a}$ denotes a metric-compatible derivative operator with torsion. The connection relating $\mathring{\nabla}_{a}$ and $\nabla_{a}$ is given by
@@
for an arbitrary one-form $v_{a}$. This connection is also known as the contorsion tensor, and it is related to torsion via
$K^{c}_{\ ab}=\frac{1}{2}\big(T^{c}_{\ ab}+T_{a\ b}^{\ c}+T_{b\ a}^{\ c}\big).$ (8)
Note that $K_{cab}=K_{[c|a|b]}$, with $K_{cab}=g_{cd}K^{d}_{\ ab}$, and $T^{c}_{\ ab}=2K^{c}_{\ [ab]}$. Hence, torsion can always be determined from contorsion, and vice versa. Accordingly, in what follows, we use these two fields interchangeably.
@@
$R_{abc}^{\quad d}\,v_{d}=\big(\nabla_{a}\nabla_{b}-\nabla_{b}\nabla_{a}+T^{d}_{\ ab}\nabla_{d}\big)v_{c},$ (9)
@@
$R_{abc}^{\quad d}=\mathring{R}_{abc}^{\quad d}-2\,\mathring{\nabla}_{[a}K^{d}_{\ b]c}+2\,K^{e}_{\ [a|c}K^{d}_{\ |b]e}.$ (10)
In addition, we can readily verify that $R_{abc}^{\quad d}=R_{[ab]c}^{\quad\ \ d}$, and, from metric compatibility, $R_{abcd}=R_{ab[cd]}$ (with $R_{abcd}=g_{de}R_{abc}^{\quad e}$). Moreover,
$R_{[abc]}^{\quad\ \ d}$ $=\nabla_{[b}T^{d}_{\ a]c}+T^{e}_{\ [ab}T^{d}_{\ c]e},$ (11)
$\nabla_{[a}R_{bc]d}^{\quad\ e}$ $=T^{f}_{\ [ab}R_{c]fd}^{\quad\ e},$ (12)
@@
$R_{ab}$ $=\mathring{R}_{ab}-2\,\mathring{\nabla}_{[a}K^{c}_{\ c]b}+2\,K^{d}_{\ [a|b}K^{c}_{\ |c]d},$ (13)
$R$ $=\mathring{R}+2\,\mathring{\nabla}_{a}K^{a\ b}_{\ b}+2\,K^{c\ \ a}_{\ [a}K^{b}_{\ b]c}.$ (14)
@@ 3.2 Classical Dynamics
$S[g^{ab},K^{c}_{\ ab},\Phi]=\frac{1}{16\pi}\int R\,\sqrt{-g}\,\mathrm{d}^{4}x+S_{\text{M}},$ (15)
where the first term has the same form as the Einstein–Hilbert action of GR, but with a torsionful Ricci scalar, and $S_{\text{M}}=\int\mathcal{L}_{\text{M}}\sqrt{-g}\,\mathrm{d}^{4}x$, with $\mathcal{L}_{\text{M}}=\mathcal{L}_{\text{M}}[g^{ab},K^{c}_{\ ab},\Phi]$. We take the variation of the action, neglecting boundary terms (as we do throughout the text), and obtain the equations of motion:
$\mathring{G}_{ab}+K^{d}_{\ (ab)}K^{c}_{\ cd}-K^{d}_{\ c(a}K^{c}_{\ b)d}$ $-K^{d\ \ c}_{\ [c}K^{e}_{\ e]d}g_{ab}$
@@
$g^{ab}K^{d}_{\ dc}-\delta^{a}_{\ c}K^{d\ b}_{\ d}+K_{c}^{\ ba}-K^{b\ a}_{\ c}$ $=16\pi\sigma_{c}^{\ ab},$ (17)
@@
$\sigma_{c}^{\ ab}$ $=-\frac{1}{\sqrt{-g}}\frac{\delta(\mathcal{L}_{\text{M}}\sqrt{-g})}{\delta K^{c}_{\ ab}}.$ (20)
We note that $\tau_{ab}=\tau_{(ab)}$ and $\sigma^{cab}=\sigma^{[c|a|b]}$, where $\sigma^{cab}=g^{cd}\sigma_{d}^{\ ab}$.
Equation (16) generalizes Einstein’s field equations, while Eq. (18) provides the equations of motion for the matter fields. Notably, Eq. (17) is algebraic, implying that the contorsion does not propagate (theories allowing for propagating torsion have been proposed; see, e.g., Ref. [34]). In fact, we can invert Eq. (17) to show that $K^{c}_{\ ab}\neq 0$ only where $\sigma_{c}^{\ ab}\neq 0$, which typically occurs inside matter.
@@
$+\xi g^{ab}[2(K^{c}_{\ cb}\Phi\mathring{\nabla}_{a}\Phi-K^{c}_{\ ab}\Phi\mathring{\nabla}_{c}\Phi)$
$+(K^{d}_{\ ab}K^{c}_{\ cd}-K^{d}_{\ cb}K^{c}_{\ ad})\Phi^{2}]\}\sqrt{-g}\text{d}^{4}x.$ (23)
@@
$+\xi[2(K^{c}_{\ c(a}\Phi\mathring{\nabla}_{b)}\Phi-K^{c}_{\ (ab)}\Phi\mathring{\nabla}_{c}\Phi)$
$+(K^{d}_{\ (ab)}K^{c}_{\ cd}-K^{d}_{\ c(a}K^{c}_{\ b)d})\Phi^{2}$
$-(K^{d\ \ c}_{\ [c}K^{e}_{\ e]d}\Phi^{2}-2K^{c\ d}_{\ d}\Phi\mathring{\nabla}_{c}\Phi)g_{ab}],$ (24)
@@
$\sigma_{c}^{\ ab}$ $=\xi(\delta^{a}_{\ c}\Phi\mathring{\nabla}^{b}\Phi-g^{ab}\Phi\mathring{\nabla}_{c}\Phi)$
$+\frac{1}{2}\xi(g^{ab}K^{d}_{\ dc}-\delta^{a}_{\ c}K^{d\ b}_{\ d}+K_{c}^{\ ba}-K^{b\ a}_{\ c})\Phi^{2}.$ (25)
@@
Another point worth noting concerns the case of vanishing contorsion. When $K^{c}_{\ ab}=0$, Eq. (3.2) reduces to
$0=\xi\big(\delta^{a}_{\ c}\Phi\mathring{\nabla}^{b}\Phi-g^{ab}\Phi\mathring{\nabla}^{c}\Phi\big),$ (26)
which implies that the scalar field must be constant. Hence, contrary to naive expectations, setting $K^{c}_{\ ab}=0$ does not reproduce GR with a generic Klein–Gordon field. To obtain this theory, the contorsion must be set to zero before varying the action. This subtlety is important in the semiclassical analysis, because, within the axiomatic framework, we work directly with the equations of motion; thus, $K^{c}_{\ ab}=0$ does not reduce to the semiclassical Klein–Gordon field propagating in a torsionless spacetime.
@@ 3.3.1 Invariance under diffeomorphisms
$\delta S_{\text{M}}=-\int\Bigg\{\frac{1}{2}\tau_{ab}\delta g^{ab}+\sigma_{c}^{\ ab}\delta K^{c}_{\ ab}-\frac{\delta\mathcal{L}_{\text{M}}}{\delta\Phi}\delta\Phi\Bigg\}\sqrt{-g}\text{d}^{4}x,$ (27)
@@
$\delta K^{c}_{\ ab}$ $=-\epsilon^{d}\mathring{\nabla}_{d}K^{c}_{\ ab}+K^{d}_{\ ab}\mathring{\nabla}_{d}\epsilon^{c}$
$-K^{c}_{\ db}\mathring{\nabla}_{a}\epsilon^{d}-K^{c}_{\ ad}\mathring{\nabla}_{b}\epsilon^{d},$ (29)
@@
$\delta S_{\text{M}}$ $=-\int\epsilon^{b}\{-\mathring{\nabla}^{a}\tau_{ab}-\sigma_{c}^{\ ad}\mathring{\nabla}_{b}K^{c}_{\ ad}-\mathring{\nabla}_{d}K^{d}_{\ ac}\sigma_{b}^{\ ac}$
$-K^{d}_{\ ac}\mathring{\nabla}_{d}\sigma_{b}^{\ ac}+\mathring{\nabla}_{a}K^{c}_{\ bd}\sigma_{c}^{\ ad}+K^{c}_{\ bd}\mathring{\nabla}_{a}\sigma_{c}^{\ ad}$
$+\mathring{\nabla}_{d}K^{c}_{\ ab}\sigma_{c}^{\ ad}+K^{c}_{\ ab}\mathring{\nabla}_{d}\sigma_{c}^{\ ad}\}\sqrt{-g}\text{d}^{4}x.$ (31)
@@
$\mathring{\nabla}^{a}\tau_{ab}$ $=(K^{c}_{\ bd}\mathring{\nabla}_{a}+K^{c}_{\ ab}\mathring{\nabla}_{d}+\mathring{\nabla}_{a}K^{c}_{\ bd}+\mathring{\nabla}_{d}K^{c}_{\ ab})\sigma_{c}^{\ ad}$
$-(K^{d}_{\ ac}\mathring{\nabla}_{d}+\mathring{\nabla}_{d}K^{d}_{\ ac})\sigma_{b}^{\ ac}-\mathring{\nabla}_{b}K^{c}_{\ ad}\sigma_{c}^{\ ad}.$ (32)

### v2

@@ 1 Introduction
Importantly, the empirical success of GR does not rule out the presence of torsion. In fact, Einstein–Cartan theory reproduces all experimentally tested predictions of GR [12]. What is more, within the framework of the Standard Model, possible signatures of torsion can be sought as new interactions involving polarized Dirac spinors [13].
@@
We organize the paper as follows. In Sec. 2, we provide a brief overview of semiclassical gravity and introduce some technical aspects. Section 3 presents the Einstein–Cartan theory, and in Sec. 4, we derive the corresponding Hadamard bi-parametrix. In Sec. 5, which constitutes the core of the paper, we introduce the axiomatic framework employed in the renormalization scheme, identify the resulting ambiguities, and review the conformal anomaly. Finally, Sec. 6 contains our conclusions. Some useful expressions are collected in Appendix A.
@@ 2 Preliminaries
We begin by discussing the left-hand side of Eq. (1). The metric curvature tensor, ${\mathring{R}_{abc}}^{\phantom{abc}d}$, is constructed using the torsion-free and metric-compatible derivative operator $\mathring{\nabla}_{a}$. These two properties mean, respectively, that
@@
The only nontrivial trace of ${\mathring{R}_{abc}}^{\phantom{abc}d}$ is $\mathring{R}_{ab}={\mathring{R}_{acb}}^{\phantom{abc}c}$, which defines the Ricci tensor and satisfies $\mathring{R}_{ab}=\mathring{R}_{ba}$. Moreover, the trace of the Ricci tensor is the Ricci scalar, $\mathring{R}=g^{ab}\mathring{R}_{ab}$. Finally, the Einstein tensor is given by
@@
As we shall see, this issue is resolved by acting with the energy–momentum operator on the Green function, $G(x,x^{\prime})$, and then taking the coincidence limit $x^{\prime}\rightarrow x$. This evaluation is divergent, and a renormalization procedure is required. To this end, Hadamard states are employed; these are states whose singular structure mirrors that of quantum fields in flat spacetime [30, 31, 32]. This allows for the subtraction of the singular part (an alternative approach is discussed in Ref. [33]). Importantly, in semiclassical GR, an axiomatic framework exists [22] that leads to a well-defined expectation value for the energy–momentum tensor.
Our goal is to study semiclassical gravity in Einstein–Cartan theory, where spacetime torsion is nontrivial. The procedure we employ closely follows that of Décanini and Folacci [32]. The most notable difference compared with GR is that the energy–momentum tensor is no longer divergence-free, a property that is repeatedly used in Ref. [32]. Instead, in Einstein–Cartan theory, its divergence satisfies a specific identity that also involves torsion and the spin–density tensor. We now turn to study some classical aspects of this theory.
@@ 3.1 Kinematics
We consider a nontrivial torsion tensor $T^{c}_{\phantom{c}ab}$, defined by
$(\nabla_{a}\nabla_{b}-\nabla_{b}\nabla_{a})f=-T^{c}_{\ ab}\nabla_{c}f,$ (6)
for any function $f$. By definition, $T^{c}_{\phantom{c}ab}=T^{c}_{\phantom{c}[ab]}$. Here, $\nabla_{a}$ denotes a metric-compatible derivative operator with torsion. The operators $\mathring{\nabla}_{a}$ and $\nabla_{a}$ are related by
@@
for an arbitrary one-form $v_{a}$, where
$K^{c}_{\phantom{c}ab}=\frac{1}{2}\left(T^{c}_{\phantom{c}ab}+T_{a\phantom{c}b}^{\phantom{a}c}+T_{b\phantom{c}a}^{\phantom{b}c}\right),$ (8)
is the contorsion tensor. Note that $K_{cab}=K_{[c|a|b]}$, with $K_{cab}=g_{cd}K^{d}_{\phantom{d}ab}$, and $T^{c}_{\phantom{c}ab}=2K^{c}_{\phantom{c}[ab]}$. Hence, torsion can always be determined from contorsion, and vice versa. Accordingly, in what follows, we use these two fields interchangeably.
@@
$R_{abc}^{\phantom{abc}d}\,v_{d}=\big(\nabla_{a}\nabla_{b}-\nabla_{b}\nabla_{a}+T^{d}_{\phantom{d}ab}\nabla_{d}\big)v_{c},$ (9)
@@
$R_{abc}^{\phantom{abc}d}=\mathring{R}_{abc}^{\phantom{abc}d}-2\,\mathring{\nabla}_{[a}K^{d}_{\phantom{d}b]c}+2\,K^{e}_{\phantom{e}[a|c}K^{d}_{\phantom{d}|b]e}.$ (10)
In addition, we can readily verify that $R_{abc}^{\phantom{abc}d}=R_{[ab]c}^{\phantom{[ab]c}d}$, and, from metric compatibility, $R_{abcd}=R_{ab[cd]}$ (with $R_{abcd}=g_{de}R_{abc}^{\phantom{abc}e}$). Moreover,
$R_{[abc]}^{\phantom{[abc]}d}$ $=\nabla_{[b}T^{d}_{\phantom{d}a]c}+T^{e}_{\phantom{e}[ab}T^{d}_{\phantom{d}c]e},$ (11)
$\nabla_{[a}R_{bc]d}^{\phantom{bc]d}e}$ $=T^{f}_{\phantom{f}[ab}R_{c]fd}^{\phantom{c]fd}e},$ (12)
@@
$R_{ab}$ $=\mathring{R}_{ab}-2\,\mathring{\nabla}_{[a}K^{c}_{\phantom{c}c]b}+2\,K^{d}_{\phantom{d}[a|b}K^{c}_{\phantom{c}|c]d},$ (13)
$R$ $=\mathring{R}+2\,\mathring{\nabla}_{a}K^{a\phantom{b}b}_{\phantom{a}b}+2\,K^{c\phantom{[a}a}_{\phantom{c}[a}K^{b}_{\phantom{b}b]c}.$ (14)
@@ 3.2 Classical Dynamics
$S[g^{ab},K^{c}_{\phantom{c}ab},\Phi]=\frac{1}{16\pi}\int R\,\sqrt{-g}\,\mathrm{d}^{4}x+S_{\text{M}},$ (15)
where the first term has the same form as the Einstein–Hilbert action of GR, but with a torsionful Ricci scalar, and $S_{\text{M}}=\int\mathcal{L}_{\text{M}}\sqrt{-g}\,\mathrm{d}^{4}x$, with $\mathcal{L}_{\text{M}}=\mathcal{L}_{\text{M}}[g^{ab},K^{c}_{\phantom{c}ab},\Phi]$. We take the variation of the action, neglecting boundary terms (as we do throughout the text), and obtain the equations of motion:
$\mathring{G}_{ab}+K^{d}_{\phantom{d}(ab)}K^{c}_{\phantom{c}cd}-K^{d}_{\phantom{d}c(a}K^{c}_{\phantom{c}b)d}$ $-K^{d\phantom{[c}c}_{\phantom{d}[c}K^{e}_{\phantom{e}e]d}g_{ab}$
@@
$g^{ab}K^{d}_{\phantom{d}dc}-\delta^{a}_{\phantom{a}c}K^{d\phantom{d}b}_{\phantom{d}d}+K_{c}^{\phantom{c}ba}-K^{b\phantom{c}a}_{\phantom{b}c}$ $=16\pi\sigma_{c}^{\phantom{c}ab},$ (17)
@@
$\sigma_{c}^{\phantom{c}ab}$ $=-\frac{1}{\sqrt{-g}}\frac{\delta(\mathcal{L}_{\text{M}}\sqrt{-g})}{\delta K^{c}_{\phantom{c}ab}}.$ (20)
We note that $\tau_{ab}=\tau_{(ab)}$ and $\sigma^{cab}=\sigma^{[c|a|b]}$, where $\sigma^{cab}=g^{cd}\sigma_{d}^{\phantom{d}ab}$.
Equation (16) generalizes Einstein’s field equations, while Eq. (18) provides the equations of motion for the matter fields. Notably, Eq. (17) is algebraic, implying that the contorsion does not propagate (theories allowing for propagating torsion have been proposed; see, e.g., Ref. [34]). In fact, we can invert Eq. (17) to show that $K^{c}_{\phantom{c}ab}\neq 0$ only where $\sigma_{c}^{\phantom{c}ab}\neq 0$, which typically occurs inside matter.
@@
$+\xi g^{ab}[2(K^{c}_{\phantom{c}cb}\Phi\mathring{\nabla}_{a}\Phi-K^{c}_{\phantom{c}ab}\Phi\mathring{\nabla}_{c}\Phi)$
$+(K^{d}_{\phantom{d}ab}K^{c}_{\phantom{c}cd}-K^{d}_{\phantom{d}cb}K^{c}_{\phantom{c}ad})\Phi^{2}]\}\sqrt{-g}\text{d}^{4}x.$ (23)
@@
$+\xi[2(K^{c}_{\phantom{c}c(a}\Phi\mathring{\nabla}_{b)}\Phi-K^{c}_{\phantom{c}(ab)}\Phi\mathring{\nabla}_{c}\Phi)$
$+(K^{d}_{\phantom{d}(ab)}K^{c}_{\phantom{c}cd}-K^{d}_{\phantom{d}c(a}K^{c}_{\phantom{c}b)d})\Phi^{2}$
$-(K^{d\phantom{[c}c}_{\phantom{d}[c}K^{e}_{\phantom{e}e]d}\Phi^{2}-2K^{c\phantom{d}d}_{\phantom{c}d}\Phi\mathring{\nabla}_{c}\Phi)g_{ab}],$ (24)
@@
$\sigma_{c}^{\phantom{c}ab}$ $=\xi(\delta^{a}_{\phantom{a}c}\Phi\mathring{\nabla}^{b}\Phi-g^{ab}\Phi\mathring{\nabla}_{c}\Phi)$
$+\frac{1}{2}\xi(g^{ab}K^{d}_{\phantom{d}dc}-\delta^{a}_{\phantom{a}c}K^{d\phantom{d}b}_{\phantom{d}d}+K_{c}^{\phantom{c}ba}-K^{b\phantom{c}a}_{\phantom{b}c})\Phi^{2}.$ (25)
@@
Another point worth noting concerns the case of vanishing contorsion. When $K^{c}_{\phantom{c}ab}=0$, Eq. (3.2) reduces to
$0=\xi\big(\delta^{a}_{\phantom{a}c}\Phi\mathring{\nabla}^{b}\Phi-g^{ab}\Phi\mathring{\nabla}_{c}\Phi\big),$ (26)
which implies that the scalar field must be constant. Hence, contrary to naive expectations, setting $K^{c}_{\phantom{c}ab}=0$ does not reproduce GR with a generic Klein–Gordon field. To obtain this theory, the contorsion must be set to zero before varying the action. This subtlety is important in the semiclassical analysis, because, within the axiomatic framework, we work directly with the equations of motion; thus, $K^{c}_{\phantom{c}ab}=0$ does not reduce to the semiclassical Klein–Gordon field propagating in a torsionless spacetime.
@@ 3.3.1 Invariance under diffeomorphisms
$\delta S_{\text{M}}=-\int\Bigg\{\frac{1}{2}\tau_{ab}\delta g^{ab}+\sigma_{c}^{\phantom{c}ab}\delta K^{c}_{\phantom{c}ab}-\frac{\delta\mathcal{L}_{\text{M}}}{\delta\Phi}\delta\Phi\Bigg\}\sqrt{-g}\text{d}^{4}x,$ (27)
@@
$\delta K^{c}_{\phantom{c}ab}$ $=-\epsilon^{d}\mathring{\nabla}_{d}K^{c}_{\phantom{c}ab}+K^{d}_{\phantom{d}ab}\mathring{\nabla}_{d}\epsilon^{c}$
$-K^{c}_{\phantom{c}db}\mathring{\nabla}_{a}\epsilon^{d}-K^{c}_{\phantom{c}ad}\mathring{\nabla}_{b}\epsilon^{d},$ (29)
@@
$\delta S_{\text{M}}$ $=-\int\epsilon^{b}\{-\mathring{\nabla}^{a}\tau_{ab}-\sigma_{c}^{\phantom{c}ad}\mathring{\nabla}_{b}K^{c}_{\phantom{c}ad}-\mathring{\nabla}_{d}K^{d}_{\phantom{d}ac}\sigma_{b}^{\phantom{b}ac}$
$-K^{d}_{\phantom{d}ac}\mathring{\nabla}_{d}\sigma_{b}^{\phantom{b}ac}+\mathring{\nabla}_{a}K^{c}_{\phantom{c}bd}\sigma_{c}^{\phantom{c}ad}+K^{c}_{\phantom{c}bd}\mathring{\nabla}_{a}\sigma_{c}^{\phantom{c}ad}$
$+\mathring{\nabla}_{d}K^{c}_{\phantom{c}ab}\sigma_{c}^{\phantom{c}ad}+K^{c}_{\phantom{c}ab}\mathring{\nabla}_{d}\sigma_{c}^{\phantom{c}ad}\}\sqrt{-g}\text{d}^{4}x.$ (31)
@@
$\mathring{\nabla}^{a}\tau_{ab}$ $=(K^{c}_{\phantom{c}bd}\mathring{\nabla}_{a}+K^{c}_{\phantom{c}ab}\mathring{\nabla}_{d}+\mathring{\nabla}_{a}K^{c}_{\phantom{c}bd}+\mathring{\nabla}_{d}K^{c}_{\phantom{c}ab})\sigma_{c}^{\phantom{c}ad}$
$-(K^{d}_{\phantom{d}ac}\mathring{\nabla}_{d}+\mathring{\nabla}_{d}K^{d}_{\phantom{d}ac})\sigma_{b}^{\phantom{b}ac}-\mathring{\nabla}_{b}K^{c}_{\phantom{c}ad}\sigma_{c}^{\phantom{c}ad}.$ (32)